A study conducted by a local cable company estimates that the number of subscribers will grow at the rate of 120+280t^7/4 new subscribers per month, t months from the start date of service. If there were 800 subscribers signed up after one month:

a) Find the function for the number of subscribers during any month t.

b) How many subscribers were signed up after 16 months?
Answer #1

Solution:

The rate of increase of the number of cable TV subscriber is \small 120+280t^{\frac{7}{4}} where \small t is the month from the start date of service.

The subscriber of Cable TV after one month is \small 800 .

a)

Integrate the given expression for cable TV with respect to \small t we get

f (t) = (120 + 280 ) dt

\small =120t+\frac{1120}{11}t^{\frac{11}{4}}+C

For  \small t=1, \: \: f\left ( 1 \right )=800

\small \therefore f\left ( 1 \right )=800=120+\frac{1120}{11}+C

\small \therefore C=800-120-\frac{1120}{11}=\frac{6360}{11}

Thus, the function for the number of subscriber during any month \small t is

\small f\left ( t \right )=120t+\frac{1120}{11}t^{\frac{11}{4}}+\frac{6360}{11}

__________________________________________

b)

For the second part put  \small t=16

\small f\left ( 16 \right )=120\times 16+\frac{1120}{11}\left ( 16 \right )^{^{\frac{11}{4}}}+\frac{6360}{11}\approx 211022

Thus , the number of subscriber signed up after \small 16 months is \small 211022 .